### v1

@@ VII.1 Density of states
As mentioned earlier, the YSR regime has two unique low energy states, the ground state with root distribution Eq.(43) and the YSR state with root distribution Eq.(45). Now, we focus on calculating the spinon density of states in both the ground state and the YSR state to determine whether the impurity is screened. The impurity density of states in the ground state (that is obtained from root distribution Eq.(43) is of the form
@@
Notice that the integrated density of states contribution due to impurity is
@@
which is positive, and hence, the impurity is screened by the bound mode. Both the one string and the impurity term contribute negatively, but there is a sole positive contribution that comes from the boundary string root. Hence, this shows that this is a single mode of screening for impurity.
@@ IX Thermodynamics in the Kondo phase
The impurity free energy (see Appendix E for details) is of the form
@@
We plot the following for various values of $m$. Notice that at $T\to 0$, the entropy $S=\ln\Omega$ is given by a non-integer $\Omega=\sqrt{2}$ which shows that the impurity exhibits non-Fermi liquid behavior. This is consistent with the fact that the spin-1 fermions in the bulk overscreens the spin-$\frac{1}{2}$ impurity in the boundary.
@@
Figure 5: Impurity entropy for $m=0.5$ (shown in dashed lines) and for $m=1$ (shown in solid lines) obtained by numerically solving Eq.(53) using Eq.(59) as the driving term. The plot slows that irrespective of the RG invariant quantity $d$, the impurity entropy is $\ln\sqrt{2}$ at $T\to 0$, and it smoothly crosses over to $\ln 2$ as $T\to\infty$.
@@ X Discussion
In this work, we studied the physics of a single spin-$\frac{1}{2}$ magnetic impurity antiferromagnetically coupled to the edge of a one-dimensional spin-$1$ superconductor modeled by an integrable Gross-Neveu like O(3) invariant model with attractive spin interaction between the fermions of opposite chirality. We find that the bulk coupling constant $g$ always flows to a strong coupling constant in the regime studied, and hence, the model generates a mass gap $2m$. The boundary impurity, however, exhibits different behavior depending on the relative strength of the bulk coupling $g$ and the Kondo coupling $J$. When $J\gg g$, the running coupling constant $J$ flows to the intermediate 2-channel Kondo fixed point where the impurity is overscreened by the multiparticle Kondo cloud. Computing the low-temperature entropy of the impurity, we show that the impurity entropy is $\frac{1}{2}\ln 2$ consistent with the two-channel Kondo problem. When $J\ll g$, the bulk superconducting order wins such that the impurity coupling flows to zero, and hence, the impurity is unscreened as if the impurity coupling were ferromagnetic. Whereas when the two couplings are of comparable strength, two novel phases arise: a) the zero mode phase, where the impurity is overscreened by the Kondo cloud while there exists a zero energy boundary excitation described by the boundary string solution of the Bethe Ansatz equations and b) the YSR phase where the boundary string solution acquires finite mass smaller than the bulk mass gap $2m$ and hence there exists a mid-gap elementary boundary that screens the impurity while the ground state hosts unscreened impurity.
@@ B.1 One particle sector
and $I^{j0}$ is the identity operator; throughout, we use the convention that subscripts denote the spaces in which the operators act non trivially. Since in one particle sector, the only interaction involves the fermion interacting with the impurity and changing its chirality, the identification of the impurity-electron S-matrix $S^{j0}$ completes the construction of the eigenstate |$\ket{k}$ which has eigenvalue $E=k$.
@@ B.2 Two particle sector
We shall now consider the two-particle sector sectors where the bulk interaction with coupling strength $g$ in the Hamiltonian Eq.(66) comes into play. As usual, we shall write the wave function as the sum of plane waves with different amplitudes in different regions that are separated by the ordering of the particles i.e.
@@
where the sum is taken over all spin and chirality configuration and $F_{a_{i}a_{j}a_{0}}^{\sigma_{i}\sigma_{j}}(x_{i},x_{j})$ is the two-particle wave function, which is written in terms of the scattering amplitudes as in different sectors depending on the ordering of the particles as
@@
Here, $A_{a_{i}a_{j}a_{0}}^{\sigma_{i}\sigma_{j}}[\mathcal{C}]$ are the amplitudes for configurations with specified spin $a_{j}$ and chirality $\sigma_{j}$ as well as the ordering of particles in the configuration space, for example $\mathcal{C}=ij$ refers to the configuration where $i^{\mathrm{th}}$ is to the left of the $j^{\mathrm{th}}$ particle and $\mathcal{C}=ji$ refers to the configuration where the particles are exchanged.
@@
Unlike the scattering matrices $S^{ij}$, the scattering matrices $W^{ij}$ are not dictated by the Hamiltonian as there is no interaction between the particles of the same chirality. Their matrices are fixed via a consistency relation called the reflection equation
@@
Having obtained these various scattering matrices, we have now completed the diagonalization of the Hamiltonian in the two particles sector.
@@ Appendix C Integrability of the model
This transfer matrix is an operator that ‘transports’ the $j^{\mathrm{th}}$ particle from one end of the system to the other and returns it to the original place, gathering $S$ matrix factors along the way as it scatters off the remaining $N-1$ particles, first as a right mover and then as a left mover once its chirality changes as it scatters off the impurity. Using the relations Eq.(85)-Eq.(87), it is easy to prove that the transfer matrices $Z_{j}$ and $Z_{k}$ for any $j,k$ commute i.e. $[Z_{j},Z_{k}]=0$ which shows that all of these transfer matrices can be diagonalized simultaneously. In order to diagonalize these transfer matrices, we use functional Bethe Ansatz method [64, 65] and fusion hierarchy [65, 66, 67, 68]. In order to proceed with this formalism, we need to embed the bare S-matrices derived previously into a continuous framework. This means determining a spectral parameter-dependent R-matrix $R(\lambda)$ and a boundary reflection matrix $K(\lambda)$ so that, for particular values of the spectral parameter $\lambda$, we can obtain all the previously derived bare S-matrices for the model.
@@ C.1 Bethe Ansatz Equations
In this section, we shall write the monodromy matrix with auxiliary space $A$ taken to be that of spin-$\frac{1}{2}$ particle. Then the monodromy matrix can be written as
@@
However, to derive the relation between these Bethe roots and the transfer matrix, we need to consider the transfer matrix defined with auxiliary space of spin 1 particle.
Writing out the monodromy matrix with auxiliary space taken to be that of spin-1 particle
@@ Appendix D Alternative derivation of the Bethe Ansatz equations
One of us studied the problem in the absence of impurity with periodic boundary conditions previously [33]. There we obtained the Bethe equations by performing the symmetric fusion at the level of Bethe equations by noticing that the spin-$1$ model can be constructed by taking the two copies of spin-$\frac{1}{2}$ models and performing symmetric fusion. Generalizing the detailed construction outlined in [33] for the case of periodic boundary conditions to our case with open boundary conditions and impurity, we shall derive the Bethe equations for our present model in this section.
@@
Now, using the fact that the eigenvalues of the bare transfer matrix $e^{2ik_{j}L}$ are related to the eigenvalues of transfer matrix Eq.(103), following the same routine described above, we arrive at the Bethe equation
@@
We shall now take 2-copies of these spin-$\frac{1}{2}$ Gross-Neveu model, one with impurity and one without impurity and write the Bethe equations as
@@
As explained in the main text, the total number of spin-$\frac{1}{2}$ particles is always even in the bulk as it is constructed from two copies of $SU(2)_{1}$ Gross-Neveu model with equal numbers of particles. When a spin-$\frac{1}{2}$ impurity is added to the boundary, the total number of spin-$\frac{1}{2}$ particle is always odd. Hence, there is a propagating spinon of spin-$\frac{1}{2}$ that is present in the ground state of the model. Moreover, it is important to notice that since the model is integrable, there is no pair production during the scattering process, and hence, it makes sense to talk about the model defined on an open line with a fixed number of particles.
@@ Appendix E Thermodynamics in the Kondo phase
$\nu_{n}(\mu)=\frac{1}{2\pi}\left[\Theta_{n}(\lambda^{(n)}_{\gamma})+\sum_{\nu}\Theta_{n}(\lambda^{(n)}_{\gamma}+\nu d)+N\Theta_{n-1}(\lambda^{(n)}_{\gamma}+\nu b)+N\Theta_{n+1}(\lambda^{(n)}_{\gamma}+\nu b)-\sum_{m,\beta}\Theta_{n,m}(\lambda^{(n)}_{\gamma}+\nu\lambda^{(m)}_{\beta})\right]$ (127)
@@
$E$ $\sum_{j}\frac{2\pi}{L}n_{j}+\sum_{n}D\int\mathrm{d}\lambda\sigma_{n}(\lambda)\left[\Theta_{n-1}\left(b-\lambda\right)+\Theta_{n+1}\left(b-\lambda\right)+\Theta_{n-1}\left(\lambda+b\right)+\Theta_{n+1}\left(\lambda+b\right)-4\pi\right].$ (133)
The free energy of the spin part can be written as
@@ Appendix F Derivation of renormalization group equations
In the main text, we obtained the expression for superconducting mass gap

### v2

@@ VII.1 Density of states
As mentioned earlier, the YSR regime has two unique low-energy states, the ground state with root distribution Eq.(44) and the YSR state with root distribution Eq.(46). Now, we focus on calculating the spinon density of states in both the ground state and the YSR state to determine whether the impurity is screened. The impurity density of states in the ground state (that is obtained from the root distribution Eq.(44) is of the form
@@
Notice that the integrated density of states contribution due to the impurity is
@@
which is positive, and hence, the impurity is screened by the bound mode. Both the one-string and the impurity term contribute negatively, but there is a sole positive contribution that comes from the boundary string root. Hence, this shows that this is a single mode of screening for impurity.
In summary, the YSR phase arises when the strength of the bulk superconducting order is further increased, making the superconducting interaction comparable to or stronger than the Kondo coupling. In this regime, the impurity is no longer screened in the ground state, but a single-particle bound mode — akin to a Yu-Shiba-Rusinov state in a BCS superconductor — can screen it in an excited state. This midgap state has a finite mass, lying below the bulk superconducting gap, and provides an alternative mechanism for impurity screening distinct from the many-body Kondo cloud seen in the overscreened phase. The presence of a single-particle bound state localized at the edge reflects the competition between Kondo physics and superconductivity, where the latter dominates at low energies, preventing full Kondo screening.
@@ IX Thermodynamics in the Kondo phase
The spin part of the impurity free energy (see Appendix E for details) is of the form
@@
We proceed to solve the coupled set of equations Eq.(54) to determine $\eta_{1}(\lambda)$ for given $T,H$ and then from it compute the free energy from Eq.(59).
@@
We plot the following for various values of $m$. Notice that at $T\to 0$, the entropy $S=\ln\Omega$ is given by a non-integer $\Omega=\sqrt{2}$, which shows that the impurity exhibits non-Fermi liquid behavior. This is consistent with the fact that the spin-1 fermions in the bulk overscreen the spin-$\frac{1}{2}$ impurity in the boundary.
@@
Figure 5: Impurity entropy for $m=0.5$ (shown in dashed lines) and for $m=1$ (shown in solid lines) obtained by numerically solving Eq.(54) using Eq.(60) as the driving term. The plot shows that irrespective of the RG invariant quantity $d$, the impurity entropy is $\ln\sqrt{2}$ at $T\to 0$, and it smoothly crosses over to $\ln 2$ as $T\to\infty$.
@@ X Discussion
We considered the behavior of a single spin-$\frac{1}{2}$ magnetic impurity coupled to the edge of a one-dimensional spin-1 superconductor. The system is described by an integrable Gross-Neveu-like $O(3)$ invariant model with attractive interactions between fermions of opposite chirality. A key feature of this model is the competition between the Kondo effect and superconductivity, leading to a dynamically generated mass gap $2m$ in the bulk and a rich phase structure for the impurity at the boundary.
Regardless of the initial coupling strength, the bulk interaction parameter $g$ always flows to strong coupling under the renormalization group (RG) flow, ensuring the formation of a superconducting gap. However, the fate of the impurity is governed by the relative strength of the Kondo coupling $J$ and the bulk interaction $g$, leading to four distinct impurity phases.
When the Kondo coupling is strong ($J\gg g$), the impurity undergoes many-body overscreening by a cloud of bulk spin-1 excitations. This results in an overscreened Kondo phase, similar to the two-channel Kondo effect, where the impurity entropy at low temperatures remains nonzero at $\frac{1}{2}\ln 2$, signaling a non-Fermi liquid fixed point. The system dynamically generates a Kondo energy scale $T_{K}$, which governs impurity screening and is distinct from the dynamically generated bulk superconducting gap $\Delta=2m$.
In contrast, when the bulk superconducting interaction dominates ($J\ll g$), the system enters a local moment (unscreened) phase, where the impurity cannot be screened at any energy scale. Here, the impurity remains decoupled from the bulk due to the strong superconducting order, behaving as a free spin-$\frac{1}{2}$ at the boundary.
For intermediate values of $J$ and $g$, two novel impurity phases emerge due to the delicate balance between Kondo screening and superconductivity. In the zero mode phase, the impurity is still overscreened by the many-body Kondo cloud, but an additional zero-energy boundary excitation appears, described by a purely imaginary Bethe root. This boundary excitation leads to an exact degeneracy between the ground state and an excited state in the thermodynamic limit, making this phase distinct from the conventional overscreened Kondo regime.
Further increasing the superconducting order relative to the Kondo coupling leads to the YSR phase, where the impurity is unscreened in the ground state but can be screened in an excited state by a single-particle bound mode localized at the edge. This bound mode, akin to a Yu-Shiba-Rusinov (YSR) state in a BCS superconductor, lies inside the superconducting gap but below the bulk excitation continuum. Unlike the Kondo phase, where the impurity is screened by a collective many-body effect, the YSR phase allows for a screening mechanism via an individual bound state, demonstrating an alternative way in which impurity physics manifests in superconducting systems.
The transitions between these phases are governed by an RG-invariant parameter $d(J,g)$, which encodes the running of both the bulk and boundary couplings. This parameter determines the stability of the impurity phases and characterizes the crossover between different regimes. The competition between Kondo physics and superconductivity thus leads to a rich boundary phase diagram, illustrating how impurity behavior is modified in strongly correlated superconducting systems.
While our current results are derived in the thermodynamic limit, it would be valuable to extend the analysis using finite-size scaling to enable direct comparisons with perturbed CFT and NRG results. Such an approach would reveal the differences in the impurity’s thermodynamic and dynamical behavior across these four phases, offering deeper insights into the interplay between impurity and bulk properties.
@@ B.1 One particle sector
and $I^{j0}$ is the identity operator; throughout, we use the convention that superscripts denote the spaces in which the operators act non-trivially. Since in one particle sector, the only interaction involves the fermion interacting with the impurity and changing its chirality, the identification of the impurity-electron S-matrix $S^{j0}$ completes the construction of the eigenstate |$\ket{k}$ which has eigenvalue $E=k$.
@@ B.2 Two particle sector
We shall now consider the two-particle sector where the bulk interaction with coupling strength $g$ in the Hamiltonian Eq.(67) comes into play. As usual, we shall write the wave function as the sum of plane waves with different amplitudes in different regions that are separated by the ordering of the particles i.e.
@@
where the sum is taken over all spin and chirality configurations and $F_{a_{i}a_{j}a_{0}}^{\sigma_{i}\sigma_{j}}(x_{i},x_{j})$ is the two-particle wave function, which is written in terms of the scattering amplitudes in different sectors depending on the ordering of the particles as
@@
Here, $A_{a_{i}a_{j}a_{0}}^{\sigma_{i}\sigma_{j}}[\mathcal{C}]$ are the amplitudes for configurations with specified spin $a_{j}$ and chirality $\sigma_{j}$ as well as the ordering of particles in the configuration space, for example, $\mathcal{C}=ij$ refers to the configuration where $i^{\mathrm{th}}$ is to the left of the $j^{\mathrm{th}}$ particle and $\mathcal{C}=ji$ refers to the configuration where the particles are exchanged.
@@
Unlike the scattering matrices $S^{ij}$, the scattering matrices $W^{ij}$ are not dictated by the Hamiltonian, as there is no interaction between the particles of the same chirality. Their matrices are fixed via a consistency relation called the reflection equation
@@
Having obtained these various scattering matrices, we have now completed the diagonalization of the Hamiltonian in the two-particle sector.
@@ Appendix C Integrability of the model
This transfer matrix is an operator that ‘transports’ the $j^{\mathrm{th}}$ particle from one end of the system to the other and returns it to the original place, gathering $S$ matrix factors along the way as it scatters off the remaining $N-1$ particles, first as a right mover and then as a left mover once its chirality changes as it scatters off the impurity. Using the relations Eq.(86)-Eq.(88), it is easy to prove that the transfer matrices $Z_{j}$ and $Z_{k}$ for any $j,k$ commute i.e. $[Z_{j},Z_{k}]=0$, which shows that all of these transfer matrices can be diagonalized simultaneously. In order to diagonalize these transfer matrices, we use the functional Bethe Ansatz method [65, 66] and fusion hierarchy [66, 67, 68, 69]. In order to proceed with this formalism, we need to embed the bare S-matrices derived previously into a continuous framework. This means determining a spectral parameter-dependent R-matrix $R(\lambda)$ and a boundary reflection matrix $K(\lambda)$ so that, for particular values of the spectral parameter $\lambda$, we can obtain all the previously derived bare S-matrices for the model.
@@ C.1 Bethe Ansatz Equations
In this section, we shall write the monodromy matrix with auxiliary space $A$ taken to be that of a spin-$\frac{1}{2}$ particle. Then the monodromy matrix can be written as
@@
However, to derive the relation between these Bethe roots and the transfer matrix, we need to consider the transfer matrix defined with the auxiliary space of a spin-1 particle.
Writing out the monodromy matrix with auxiliary space taken to be that of a spin-1 particle
@@ Appendix D Alternative derivation of the Bethe Ansatz equations
One of us studied the problem in the absence of impurity with periodic boundary conditions previously [33]. There, the Bethe equations were obtained by performing the symmetric fusion at the level of the Bethe equations by noticing that the spin-$1$ model can be constructed by taking the two copies of spin-$\frac{1}{2}$ models and performing symmetric fusion. Generalizing the detailed construction outlined in [33] for the case of periodic boundary conditions to our case with open boundary conditions and impurity, we shall derive the Bethe equations for our present model in this section.
@@
$\left(\frac{\lambda_{j}+b+\frac{i}{2}}{\lambda_{j}+b-\frac{i}{2}}\right)^{N}\left(\frac{\lambda_{j}-b+\frac{i}{2}}{\lambda_{j}-b-\frac{i}{2}}\right)^{N}\frac{\lambda_{j}+d+\frac{i}{2}}{\lambda_{j}+d-\frac{i}{2}}\frac{\lambda_{j}-d+\frac{i}{2}}{\lambda_{j}-d-\frac{i}{2}}=\prod_{\ell\neq j}^{M}\frac{\lambda_{j}-\lambda_{\ell}+i}{\lambda_{j}-\lambda_{\ell}-i}\frac{\lambda_{j}+\lambda_{\ell}+i}{\lambda_{j}+\lambda_{\ell}-i}.$ (117)
Now, using the fact that the eigenvalues of the bare transfer matrix $e^{2ik_{j}L}$ are related to the eigenvalues of the transfer matrix Eq.(104), following the same routine described above, we arrive at the Bethe equation
@@
We shall now take 2 copies of these spin-$\frac{1}{2}$ Gross-Neveu model, one with impurity and one without impurity, and write the Bethe equations as
@@
and the energy equation as
$e^{-2ik_{j}L}=\prod_{\alpha=1}^{2M}\left(\frac{\mu_{j}+\lambda_{\alpha}+\frac{i}{2}}{\mu_{j}+\lambda_{\alpha}-\frac{i}{2}}\right)\left(\frac{\mu_{j}-\lambda_{\alpha}+\frac{i}{2}}{\mu_{j}-\lambda_{\alpha}-\frac{i}{2}}\right).$ (121)
@@
As explained in the main text, the total number of spin-$\frac{1}{2}$ particles is always even in the bulk as it is constructed from two copies of $SU(2)_{1}$ Gross-Neveu model with equal numbers of particles. When a spin-$\frac{1}{2}$ impurity is added to the boundary, the total number of spin-$\frac{1}{2}$ particles is always odd. Hence, there is a propagating spinon of spin-$\frac{1}{2}$ that is present in the ground state of the model. Moreover, it is important to notice that since the model is integrable, there is no pair production during the scattering process, and hence, it makes sense to talk about the model defined on an open line with a fixed number of particles.
@@ Appendix E Thermodynamics in the Kondo phase
$\nu_{n}(\lambda)=\frac{1}{2\pi}\left[\Theta_{n}(\lambda^{(n)}_{\gamma})+\sum_{\nu}\Theta_{n}(\lambda^{(n)}_{\gamma}+\nu d)+N\Theta_{n-1}(\lambda^{(n)}_{\gamma}+\nu b)+N\Theta_{n+1}(\lambda^{(n)}_{\gamma}+\nu b)-\sum_{m,\beta}\Theta_{n,m}(\lambda^{(n)}_{\gamma}+\nu\lambda^{(m)}_{\beta})\right]$ (129)
@@
$E=$ $\sum_{j}\frac{2\pi}{L}n_{j}+\sum_{n}D\int\mathrm{d}\lambda\sigma_{n}(\lambda)\left[\Theta_{n-1}\left(b-\lambda\right)+\Theta_{n+1}\left(b-\lambda\right)+\Theta_{n-1}\left(\lambda+b\right)+\Theta_{n+1}\left(\lambda+b\right)-4\pi\right].$ (135)
Here, the first term is the charge energy
$E^{(c)}(\{n_{j}\})=\frac{2\pi}{L}\sum_{j=1}^{N^{e}}n_{j},$
such that the charge partition function becomes
$Z^{(c)}=\sum_{\{n_{j}\},\,n_{j}\geq-N^{e}}\exp\left[-\frac{1}{T}\sum_{j=1}^{N^{e}}\frac{2\pi}{L}n_{j}\right].$
This describes the thermodynamics of $N$ noninteracting spinless fermions with linear kinetic energy. In the limit $D\to\infty$, it leads to the free energy
[…]
This free energy corresponds to half the free energy of a noninteracting electron gas at zero magnetic field.
The free energy of the spin part in the presence of the magnetic field $H$ can be written as
@@ Appendix F Derivation of renormalization group equations
In the main text, we obtained the expression for the superconducting mass gap